A six-sided die is rolled three times independently. Which is more likely: a sum of 11 or a sum of 12?


1
Expert's answer
2022-01-18T16:38:15-0500

The total number of outcomes when a die is rolled 3 times is "6^3=216" outcomes


A sum of 11 is obtained with the following six combinations.

"(6,4,1),(6,3,2),(5,5,1),(5,4,2),(5,3,3),(4,4,3)".

A combination of "3" distinct numbers corresponds to "^3P_1=3\times2=6" permutations. Each combination of 3, two of which are equal corresponds to "^3P_2=3" permutations

Therefore, the combinations "(6,4,1),(6,3,2),(5,4,2)" have six permutations each while the combinations "(5,5,1),(5,3,3),(4,4,3)" have three combinations each.

Thus, total number of combinations with sum equal to 11 is 6+6+6+3+3+3=27.

The probability of getting a combination with sum equal to 11 is "{27\over216}".


A sum of 12 is obtained with the following 6 combinations.

"(4,4,4),(6,4,2),(6,3,3),(5,4,3),(6,5,1),(5,5,2)"

A combination of "3" distinct numbers corresponds to "^3P_1=3\times2=6" permutations. Each combination of 3, two of which are equal corresponds to "^3P_2=3" permutations and a combination of 3, three of which are equal corresponds to "^3P_0=1" permutations.

Therefore, the combinations, "(6,4,2),(5,4,3),(6,5,1)" have six combinations each. The combinations "(6,3,3),(5,5,2)" have three combinations each and the combinations "(4,4,4)" has one permutation.

Therefore, the total number of combinations with sum equal to 12 is 6+6+6+3+3+1=25. The probability of getting a combination with sum equal to 12 is "{25\over216}".

Of the two events, the more likely event is the one with a higher probability of occurring. Therefore, the more likely event is a sum of 11 with a probability of "27\over216" .